THE MILLIMETER/SUBMILLIMETER SPECTRUM OF TiF^{+} (X^{3}Φ_{r})

\underline{D. T. HALFEN} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of the molecular ion TiF^{+} (X^{3}Φ_{r}) has been measured using a combination of millimeterwave direct absorption methods and velocity modulation techniques. TiF^{+} was created in an AC glow discharge using a mixture of gas-phase TiCl_{4}, 10% F_{2} in helium, and argon. In contrast to TiCl^{+}, the spectra of TiF^{+} were weaker, a probable result of the synthesis method. Ten rotational transitions were measured in the frequency range of 327 to 542 GHz. All three spin-orbit components were observed, which were additionally split into doublets due to fluorine hyperfine interactions. The fine structure exhibited a regular case (a) pattern with minimal second-order spin-orbit effects. These data were fit with a case (a) Hamiltonian and rotational, fine structure, and hyperfine constants were determined. The calculated bond length of TiF^{+} was found to be shorter than that of TiF, which is thought to be due to an increase in the effective nuclear charge of titanium.